OpenStudy (anonymous):

Which describes the locus of points 4 units from (5, 1) in the coordinate plane?

jimthompson5910 (jim_thompson5910):

center: (5,1) ---> h = 5, k = 1 radius: 4 ----> r = 4 (x-h)^2+(y-k)^2=r^2 (x-5)^2+(y-1)^2=4^2 (x-5)^2+(y-1)^2=16 and that's all there is to it